OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the graphing method to solve the following system of equations. (3 points) Part 2: Provide the solution to the system. (2 points) Part 3: Explain what the solution represents on the graph. (1 point) 3x – 6y = 12 9x + 2y = –24

OpenStudy (anonymous):

@campbell_st Can you help me with this one please?

OpenStudy (anonymous):

I need help on all the Parts..

OpenStudy (campbell_st):

rewrite the 1st equation with 6y as the subject 6y = 12 - 3x divide the equation by 3 2y = 4 - x substitute into the 2nd equation 9x +( 4 -x) = -24 solve for x the solution will be the point of intersection of the 2 lines

OpenStudy (anonymous):

what would the solutions be? @campbell_st

OpenStudy (campbell_st):

well you need to solve 8x + 4 = -24 to find x

OpenStudy (anonymous):

I got x = 7/2

OpenStudy (anonymous):

I don't think that is right though

OpenStudy (anonymous):

nvm i got it

OpenStudy (campbell_st):

actually x = -7/2

OpenStudy (campbell_st):

and it is the correct value for the equations in your question